AB INITIO PREDICTION OF THE STRUCTURE AND ROTATIONAL CONSTANTS OF GROUND STATE METHYLAMINE, CH_{3}NH_{2}

\underline{D. E. WOON}, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL 61801.

Accurate theoretical predictions can provide critical information for high resolution rotational spectroscopy, including both structural parameters and potential energy surfaces associated with hindered internal motions. With recent experimental data available [V. V. Ilyushin et al., J. Mol. Spectrosc. 229, 170 (2005]), methylamine is an ideal test case for benchmarking the computational demands for asymmetric top molecules with two internal motions. Several issues will be addressed: (1) What level of theory and basis set is required to accurately describe the equilibrium structure? (2) What procedure should be used to transform the equilibrium rotational constants to reflect zero-point motion? (3) What is required to obtain accurate potentials for the internal rotation and inversion of NH_{2} with respect to the CH_{3} group? Various levels of theory – B3LYP, MP2, and RCCSD(T) – were explored in conjunction with valence and core-valence correlation consistent basis sets.